OpenStudy (anonymous):

Find the geometric mean of the pair of numbers. 242 and 8 A. 44 B. 49 C. 1872 D. 54

OpenStudy (anonymous):

can someone plas help on mine

OpenStudy (anonymous):

(242*8)^1/2

hartnn (hartnn):

Geometric mean of x and y is just \(\sqrt {xy}\)

OpenStudy (anonymous):

so how would i find my answer using that

hartnn (hartnn):

\(\large \sqrt {242 \times 8} =... \)

OpenStudy (anonymous):

So it would be 44?

hartnn (hartnn):

YES! good :)

OpenStudy (anonymous):

Yay Thanks!! :D

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Can you help me with another question i have??

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

Ok hold on ima post it seperate